Assuming each solution to be 0.10 M , rank the following aqueous solutions in order of decreasing pH.

Rank the solutions from the highest to lowest pH. To rank items as equivalent, overlap them

N2H2

Ba(OH)2

HOCL

NAOH

HCL
Answer #1
Concepts and reason

The concept used is to arrange the aqueous solution of the given salts in the order of decreasing pH{\rm{pH}} .

Fundamentals

pHpH is the measure of hydrogen ion concentration. Lower the pHpH , higher is the hydrogen ion concentration and lower is the hydroxide ion concentration.

Higher the pHpH , lower is the hydrogen ion concentration and higher is the hydroxide ion concentration.

The hydrogen ion concentration is calculated as follows:

[H+]=10pH\left[ {{{\rm{H}}^ + }} \right] = {10^{ - pH}}

The hydroxide ion concentration pOHpOH and pHpH are related as follows:

[H+][OH]=1.0×1014\left[ {{{\rm{H}}^ + }} \right]\left[ {{\rm{O}}{{\rm{H}}^ - }} \right] = 1.0 \times {10^{ - 14}}

If the pH<7{\rm{pH}} < 7 , the solution is acidic.

If the pH>7{\rm{pH}} > 7 , the solution is basic.

If the pH=7{\rm{pH}} = 7 , the solution is neutral.

The dissociation of acid produces hydronium ions. The reaction is as follows:

A+H2OA(aq)+H3O+(aq){\rm{A}} + {{\rm{H}}_2}{\rm{O}}\longrightarrow{{}}{{\rm{A}}^ - }\left( {aq} \right) + {{\rm{H}}_3}{{\rm{O}}^ + }\left( {aq} \right)

The dissociation of base produces hydroxide ions. The reaction is as follows:

B+H2OBH+(aq)+OH(aq){\rm{B}} + {{\rm{H}}_2}{\rm{O}}\longrightarrow{{}}{\rm{B}}{{\rm{H}}^ + }\left( {aq} \right) + {\rm{O}}{{\rm{H}}^ - }\left( {aq} \right)

Among the given aqueous solutions, HCl{\rm{HCl}} and HOCl{\rm{HOCl}} are the acids.

HCl{\rm{HCl}} is a strong acid.

[HCl]=0.1M\left[ {{\rm{HCl}}} \right] = 0.1{\rm{ M}}

On dissociation, it produces one mole of H+{{\rm{H}}^ + } .

So,

[H+]=1×0.1M=0.1M\begin{array}{l}\\\left[ {{{\rm{H}}^ + }} \right] = 1 \times 0.1{\rm{ M}}\\\\{\rm{ = 0}}{\rm{.1 M}}\\\end{array}

Since, the hydronium concentration is high, its pH{\rm{pH}} is small.

HOCl{\rm{HOCl}} is a weak acid.

Its dissociation is not complete.

So,

[H+]<0.1M\left[ {{{\rm{H}}^ + }} \right] < {\rm{0}}{\rm{.1 M}}

Since, the concentration is small, its pH{\rm{pH}} is small and more than that of HCl{\rm{HCl}} .

Among the given aqueous solutions, Ba(OH)2{\rm{Ba}}{\left( {{\rm{OH}}} \right)_2} , NaOH{\rm{NaOH}} and N2H2{{\rm{N}}_2}{{\rm{H}}_2} are the bases.

Ba(OH)2{\rm{Ba}}{\left( {{\rm{OH}}} \right)_2} is a strong base.

[Ba(OH)2]=0.1M\left[ {{\rm{Ba}}{{\left( {{\rm{OH}}} \right)}_2}} \right] = 0.1{\rm{ M}}

On dissociation, it produces two moles of OH{\rm{O}}{{\rm{H}}^ - } .

So,

[OH]=2×0.1M=0.2M\begin{array}{l}\\\left[ {{\rm{O}}{{\rm{H}}^ - }} \right] = 2 \times 0.1{\rm{ M}}\\\\{\rm{ = 0}}{\rm{.2 M}}\\\end{array}

Since, the concentration is high, its pH{\rm{pH}} is large.

NaOH{\rm{NaOH}} is a strong base.

[NaOH]=0.1M\left[ {{\rm{NaOH}}} \right] = 0.1{\rm{ M}}

On dissociation, it produces one mole of OH{\rm{O}}{{\rm{H}}^ - } .

So,

[OH]=1×0.1M=0.1M\begin{array}{l}\\\left[ {{\rm{O}}{{\rm{H}}^ - }} \right] = 1 \times 0.1{\rm{ M}}\\\\{\rm{ = 0}}{\rm{.1 M}}\\\end{array}

Since, the concentration is high, its pH{\rm{pH}} is large but less than that of Ba(OH)2{\rm{Ba}}{\left( {{\rm{OH}}} \right)_2} .

N2H2{{\rm{N}}_2}{{\rm{H}}_2} is a weak base.

Its dissociation is not complete.

So,

[OH]<0.1M\left[ {{\rm{O}}{{\rm{H}}^ - }} \right] < {\rm{0}}{\rm{.1 M}}

Since, the concentration is small, its pH{\rm{pH}} is small and less than that of NaOH{\rm{NaOH}} .

The order of the decreasing pH{\rm{pH}} is as follows:

Ba(OH)2>NaOH>N2H2>HOCl>HCl{\rm{Ba}}{\left( {{\rm{OH}}} \right)_2} > {\rm{NaOH}} > {{\rm{N}}_2}{{\rm{H}}_2} > {\rm{HOCl}} > {\rm{HCl}}

Ans:

Therefore, the order of the decreasing pH{\rm{pH}} is as follows:

Ba(OH)2>NaOH>N2H2>HOCl>HCl{\rm{Ba}}{\left( {{\rm{OH}}} \right)_2} > {\rm{NaOH}} > {{\rm{N}}_2}{{\rm{H}}_2} > {\rm{HOCl}} > {\rm{HCl}}